OpenStudy (anonymous):

The number of visitors to a park is expected to follow the function v(x) = 5(x − 1), where x is the number of days since opening. On the first day, there will be a ceremony with 20 people in attendance. What is the function that shows total visitors, including the ceremony? c(x) = 20 + 5(x − 1) c(x) = 20 − 5(x − 1) c(x) = 5(x − 1) − 20 c(x) = 100x − 100

OpenStudy (jack1):

which do u think it is?

OpenStudy (anonymous):

i'm thinking a

OpenStudy (jack1):

i agree regular day: visitors = 5(x+1) on ceremony day = 20 extra visitors visitors = 5(x+1) + 20 = A